# Leveraging Large Language Models for Relevance Judgments in Legal Case Retrieval

### v1

@@ Abstract.
Collecting relevant judgments for legal case retrieval is a challenging and time-consuming task. Accurately judging the relevance between two legal cases requires a considerable effort to read the lengthy text and a high level of domain expertise to extract Legal Facts and make juridical judgments. With the advent of advanced large language models, some recent studies have suggested that it is promising to use LLMs for relevance judgment. Nonetheless, the method of employing a general large language model for reliable relevance judgments in legal case retrieval is yet to be thoroughly explored. To fill this research gap, we devise a novel few-shot workflow tailored to the relevant judgment of legal cases. The proposed workflow breaks down the annotation process into a series of stages, imitating the process employed by human annotators and enabling a flexible integration of expert reasoning to enhance the accuracy of relevance judgments. By comparing the relevance judgments of LLMs and human experts, we empirically show that we can obtain reliable relevance judgments with the proposed workflow. Furthermore, we demonstrate the capacity to augment existing legal case retrieval models through the synthesis of data generated by the large language model.
Large Language Models, Data Annotation, Legal Case Retrieval
copyright: acmcopyrightjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYprice: 15.00isbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Specialized information retrieval
@@ 1. Introduction
Many practitioners still rely on keyword-based retrieval systems when searching for similar cases. However, this approach is cumbersome, has a steep learning curve, and offers low efficiency and accuracy. The researchers have made extensive attempts at the LCR task. BERT-PLI(Shao et al., 2020) models the relevance by aggregating paragraph-level interactions of case pairs. SAILER(Li et al., 2023) aims to maximize the utilization of information in annotated data. Specifically, it utilizes an encoder-decoder architecture for pre-training, which encodes the facts into vectors and uses a decoder to re-construct the vectors to the original decisions and reasoning. In addition, Wei et al. (2022) propose IOT-Match, which learns to find rationales from paired legal cases based on semantics and legal elements of sentences, with the assistance of numerous expert-annotated interpretable relevance labels. Despite the promising results of these models on specific test sets, they still rely heavily on extensively high-quality annotated data. Owing to the scarcity of annotated data, a notable concern arises regarding the limited generalizability of these models across diverse data.
Figure 1. An example of a challenge of legal relevance annotation.
@@
These concerns highlight the role of a more reliable and valid method to collect the legal relevance annotations, which is vital not only for the evaluation of models but also for the training of more sophisticated models. However, to judge relevance, one necessitates substantial domain expertise and a significant effort to comprehend lengthy texts (Shao et al., 2023). This means that manual annotations require much more complex preliminary instructions for annotators, leading to higher learning effort, greater demands on their abilities, and difficulties in expanding the scale of annotated data. In particular, the major challenges in legal cases relevance annotation are threefold: 1. Expertise-intensive: The requirement for high-level domain knowledge for accurate annotation, including the intrinsic connection between the factual scenario and the corresponding judicial interpretation, and the interplay of objective conduct and subjective attitudes in shaping the ultimate legal conclusion. 2. Lengthy-text: The query and candidates often contain thousands of words. 3. Nuance-sensitive: As shown in Figure 1, besides a comprehensive understanding of the overall context, sensitivity to nuances in context and terms is equally important. Because such minor distinctions can significantly impact judicial interpretations and decisions. E.g., ”homicide” vs ”manslaughter”, ”minor injury” vs ”minor assault”, and ”robbery” and ”Theft”. (Please refer to Section  3.1 Relevance part for further details of legal relevance).
@@
Existing work(Thomas et al., 2023; Faggioli et al., 2023) has shown that the annotation capabilities of large language models (LLMs) in general NLP tasks can rival those of crowd workers. Most efforts often involved general relevance or direct relevance signals, like question-answering, and text classification(Thalken et al., 2023; Gilardi et al., 2023). Since relevance labels in legal cases are far more complicated as mentioned earlier, our major research question is whether LLM can perform relevance annotation in the legal field. Although researchers have attempted to build legal domain-specific LLMs such as Chatlaw(Cui et al., 2023) and LawGPT(Liu Hongcheng and Yuhao, 2023), they are primarily fine-tuned for general legal Q&A tasks; moreover, due to limitations in parameters and training data volume, their understanding and reasoning capabilities are inferior to general LLMs like GPT-3.5, making it hard to employ them for processing long-text content such as legal case annotation.
@@
In this work, we will focus on developing an automatic relevance annotation method for legal case retrieval. Instead of utilizing or building specialized LLMs, which carry many serious constraints in cost and capabilities (Please refer to the next section for details), we employ one of the most powerful general LLMs, GPT-3.5, for domain annotation. To take advantage of the understanding and reasoning capabilities of the general LLM, we instruct it step-by-step(Wei et al., 2022) with minimal expert guidance. As a result, our approach can activate domain-specific knowledge within the LLM and then it can annotate relevance to align with the criteria used by experts when making judgments. Specifically, we address the challenge of Lengthy-text by decomposing the annotation process into sub-progresses, including distinct views in facts extractions and relevance annotations. In each sub-process, we elaborate expert’s reasoning process as instructions to ease the Expertise-Intensive challenge and mitigate the Nuance-Sensitivity issue. In addition, we employ adaptive retrieval of effective demonstrations to further tackle the challenge of Expertise-Intensive. Given a group of unlabeled legal cases, we also designed a strategy for efficiently collecting possible positive case pairs. Through empirical experiments, our method demonstrates a high level of consistency with human labels. Furthermore, when applying the generated synthetic data for downstream model fine-tuning, we observed a significant improvement in performance on case retrieval tasks, indirectly validating the effectiveness of our approach.
@@
  • We design an automatic workflow to leverage a general LLM to make relevance judgments for legal cases. To the best of our knowledge, this is the first attempt at automated legal case relevance annotation.
  • We evaluate and analyze the quality of automatic relevance annotations by comparing them with expert annotations. Empirical experiments demonstrate that our approach can achieve high consistency with expert annotations.
  • We use the proposed method to generate a synthetic dataset, which can be used to further fine-tune legal case retrieval models, resulting in a significant performance improvement.
@@ 2.1. Pre-trained Language Models (PLMs) in Legal case retrieval
Pre-trained Language Models (PLMs) like BERT(Devlin et al., 2018), GPT-2(Lagler et al., 2013), and RoBERTa(Liu et al., 2019) have set new standards in a variety of NLP tasks. These models are trained on vast corpora to learn linguistic patterns and can be fine-tuned on specific tasks. The strength of PLMs lies in their ability to understand context, making them effective for tasks ranging from sentiment analysis to machine translation. However, in the legal domain such as legal case retrieval, general PLMs still struggle with the complex nuances and intensive background domain knowledge of legal jargon and may produce biased or asymmetric predictions. Additionally, legal documents typically consist of thousands of tokens, far exceeding the length that mainstream PLMs can handle(Xiao et al., 2021). Researchers are dedicated to designing specialized models(Shao et al., 2020; Li et al., 2023; Chalkidis et al., 2020; Xiao et al., 2021) tailored to legal case matching, underscoring the limitations of generic PLMs in accurately interpreting and classifying extensive legal content.
@@
Large-scale Language Models (LLMs) are an extension of the PLM concept but on a much grander scale, such as GPT4(OpenAI, 2023) and LLaMa(Touvron et al., 2023). With billions or even trillions of parameters, LLMs are designed to understand and generate human-like text with little to no fine-tuning required for specific tasks, showing significant potential in various fields, including law.
@@
Although researchers have attempted to train domain-specific LLMs for the legal field, such as Chatlaw(Cui et al., 2023), LawGPT(Liu Hongcheng and Yuhao, 2023), and others, there are still several drawbacks in directly training a dedicated large language model for this domain: 1. The process requires a substantial amount of high-quality data and expensive computational resources. Legal texts can be vast and diverse, necessitating extensive training data to ensure the model’s proficiency in understanding the nuances of legal language and reasoning. 2. The legal domain encompasses numerous downstream tasks, each with its own specific requirements and nuances. As a result, domain-specific large-scale language models lack generalizability and reliability across various downstream tasks. Adapting these models to each specific task requires fine-tuning with task-specific data, leading to considerable efforts for collecting specialized training data for each task. 3. Compared to general large models like GPT-3.5, these domain-specific large models suffer from insufficient training scale and data volume, leading to weaker model memorization, comprehension, and reasoning capabilities. Additionally, most of these domain-specific models have mainly undergone fine-tuning for legal question-answering tasks.
@@
In this article, we present a lightweight approach that enables general large-scale language models for legal relevance annotation.
2.2. Prompt Engineering in Data Augmentation
@@
Prompt-based data augmentation is a promising avenue, especially for tasks where data collection is challenging. The ”PromDA” model, introduced by (Wang et al., 2022a), leverages soft prompts in pre-trained language models to generate high-quality synthetic data for low-resource natural language understanding tasks. Similarly, (Wang et al., 2022b) proposed a method that combines semantic data augmentation with distance metric learning for domain generalization. (Meyer et al., 2022) explored the feasibility of GPT-3 generated synthetic data for training conversational AI classifiers, revealing that while synthetic data offers advantages, it doesn’t surpass real user data in performance. In addition, (Saad-Falcon et al., 2023) develop a method for using large language models (LLMs) to generate large numbers of synthetic queries cheaply, which generates a small number of synthetic queries using an expensive LLM and then uses a much less expensive one to create large numbers of synthetic queries that are used to fine-tune re-rankers. These studies underscore the growing importance and potential of prompt-based data augmentation in enhancing model performance, especially in scenarios with limited data resources. While it offers a novel way to leverage the capabilities of large-scale language models, care must be taken to ensure that the generated data aligns well with real-world scenarios and doesn’t introduce biases or errors. As with any data augmentation technique, it’s crucial to validate its effectiveness empirically on the target task. Therefore, in this work, we will evaluate our automatic annotation approach from two levels: the consistency with human labeling and the empirical effectiveness of data-augmented models on the LCR task.
@@ 3. Methodology
In this section, we will first provide a brief overview of the background knowledge, including an introduction to the forms of Chinese legal cases and the factors that experts consider when determining the relevance of legal cases. Then, we will provide a detailed explanation of each step in the proposed automated annotation workflow. Subsequently, we address some questions that readers may be concerned about. Finally, we introduce an application of the proposed annotation method: the construction of a synthetic dataset for training LCR models.
Figure 2. The proposed workflow of data annotation for legal case retrieval task.
@@ 3.1. Preliminary
Legal case documents diverge significantly from the general domain content. Their distinctive feature lies in their structured composition, providing both clarity and complexity. This structure is pivotal for the legal profession, as it captures the essence and progression of a case, from presenting the parties involved to the court’s final decision, which basically includes five parts: Procedure: This is the foundational section of any legal case document. It introduces essential information about the parties involved and outlines the procedural posture. Fact: Beyond mere introductions, this section delves into the core contentions of the parties. Here, one finds descriptions of arguments, evidence presented, and pivotal events forming the crux of the legal dispute. This is the only information that a pending (query) case has. Judgment process: It explains the selection of specific rules and their application to the case facts. Verdict: This section offers the court’s final conclusive response to the case.
We will evaluate our annotation method on LeCaRD(Ma et al., 2021), the Chinese legal case retrieval dataset, of which the data format is shown in Figure 3.
Relevance: According to LeCaRD, the determination of case relevance primarily depends on the key facts of the case, which are composed of ”Material Facts (MF)” and ”Legal Facts (LF)”. Legal Facts are the legal evaluations of Material Facts. For example, in a hypothetical case:
Fact: A had persistent conflicts with B. Consequently, A found a pretext to provoke B and then assaulted him. B’s injuries were classified as ’Level 2 minor injury.
@@
To determine whether different cases are related, one should deep into the Material Facts and the Legal Facts. The Material Facts of a case include the identity of multiple factors such as the defendant and the victim, the cause, process, outcome, time, and location of the crime committed by the defendant. Judging the relevance of MF involves a ”factual level” assessment of the relatedness of different cases, focusing on whether there is ”identity” between multiple factors. Judging the relevance of LF includes the subjective motive of the defendant, whether the victim is at fault, and whether there is a causal relationship between the criminal act and the damage caused. This involves an adaptive ”legal level” assessment of in different legal aspects, focusing on whether there is ”homogeneity” between two cases. In practice, the judgment should follow the order of ”Material Facts first, Legal Facts second.” Both are necessary and sufficient conditions for the assessment of case-relevance.
@@
3.2. Automated Relevance Annotation for Legal Cases
@@
We have meticulously designed a prompt workflow for legal case retrieval, which can be segmented into four steps: 1) Preliminary factual analysis by legal professionals, 2) Adaptive Demo-Matching (ADM), 3) Fact Extraction (FE), and 4) Fact Annotation (FA). Figure 2 shows the overall annotation workflow. In the subsequent sections, we will delve into each of these components in detail.
@@
Preliminary factual analysis by legal professionals: While the relevance annotations of the original data were made based on the criteria mentioned in Section 3.1, the annotated data only contain relevance scores without explicit matching labels, preventing the understanding of the exact facts that determine the similarity between two cases. Accurate and direct indications of what the ’Material Fact’ in the current case is, its connection to the details in the text, followed by additional legal knowledge and reasoning logic used for identifying such connections, can trigger LLMs to utilize their inherent knowledge and comprehension abilities to mimic human experts’ cognitive processes. Such an approach is deemed necessary because we observed that without detailed relevance indications, the model’s judgment could deviate. For instance, if case 1 involves person A stealing a car and case 2 involves person B stealing a wallet, the LLM sometimes perceives a car and a wallet as dissimilar items, focusing on trivial facts, thereby simply judging the two cases as irrelevant. Thus We consider such legal reasoning order as a logical chain, prompting the LLM to make judgments consistent with those of the professionals. Figure 4 and Figure 5 shows examples of experts’ demonstrations for extractions of MF and LF.
Figure 4. The demonstration of extractions of MF for few-shot prompting. (Translation of Chinese texts.)
Figure 5. The demonstration of extractions of LF for few-shot prompting. (Translation of Chinese texts.)
Figure 6. The demonstrations of MF and LF annotation for few-shot prompting. (Translation of Chinese texts.)
@@
It’s worth mentioning that, as stated in section 3.1, in practice, the judgment should follow the order of ’Material Facts first, Legal Facts second’. The extraction of Legal Facts, therefore, involves a multi-dimensional analysis built upon Material Facts, encompassing considerations of the defendant’s subjective motive, the potential fault of the victim, and the existence of a causal relationship between the criminal act and the resulting harm. Then during the annotation phase, it is crucial to perform adaptive relevance assessments tailored to different case pairs, based on the connection between Material Facts (MF) and Legal Facts (LF). We incorporate carefully designed expert reasoning logic of analyzing relevance to instruct the LLm in adhering to the criterion of experts. Figure 6 shows an example of demonstrations for the fact annotation.
Adaptive Demo-Matching (ADM): When applying few-shot prompting to LLMs, we adaptively fit appropriate demonstrations for input cases by retrieval in each prompting step from the small-scale pre-built expertise demonstration set. By doing so, as opposed to random selection, we can minimize the disparity between the demonstration and the input cases, facilitating the LLM to better follow legal experts. In this work, we employ a simple but effective method, BM25(Robertson and Walker, 1994), to rapidly retrieve demonstrations, which could be improved by more sophisticated models in the future.
@@
Few-shot annotation (FA): As illustrated in Figure 8, we present an framework of FA prompting. Similarly, we separately annotate MF and LF sequentially. Given as input the facts of a pair of cases to be labeled, the prompt contains relevant as well as irrelevant demonstrations. Ultimately, the LLMs are instructed to assess the relevance of the current input pair, drawing upon the expert’s rationale. The calculation of the final relevance score follows the method of LeCaRD: score 1 for MF relevance, and score 2 for LF relevance, with the total relevance score calculated as the sum of MF and LF relevance scores.
Figure 8. The example of few-shot annotation FA. (Translation of Chinese texts.)

### v2

@@ Abstract.
Determining which legal cases are relevant to a given query involves navigating lengthy texts and applying nuanced legal reasoning. Traditionally, this task has demanded significant time and domain expertise to identify key Legal Facts and reach sound juridical conclusions. In addition, existing data with legal case similarities often lack interpretability, making it difficult to understand the rationale behind relevance judgments. With the growing capabilities of large language models (LLMs), researchers have begun investigating their potential in this domain. Nonetheless, the method of employing a general large language model for reliable relevance judgments in legal case retrieval remains largely unexplored. To address this gap in research, we propose a novel few-shot approach where LLMs assist in generating expert-aligned interpretable relevance judgments. The proposed approach decomposes the judgment process into several stages, mimicking the workflow of human annotators and allowing for the flexible incorporation of expert reasoning to improve the accuracy of relevance judgments. Importantly, it also ensures interpretable data labeling, providing transparency and clarity in the relevance assessment process. Through a comparison of relevance judgments made by LLMs and human experts, we empirically demonstrate that the proposed approach can yield reliable and valid relevance assessments. Furthermore, we demonstrate that with minimal expert supervision, our approach enables a large language model to acquire case analysis expertise and subsequently transfers this ability to a smaller model via annotation-based knowledge distillation.
Relevance Judgments, Legal Case Retrieval, Synthetic Data, Large Language Models
ccs: Information systems Specialized information retrieval
@@ 1. Introduction
Many practitioners still rely on keyword-based retrieval systems when searching for similar cases. However, this approach is cumbersome, has a steep learning curve, and offers low efficiency and accuracy. The researchers have made extensive attempts at the LCR task. BERT-PLI(Shao et al., 2020) models the relevance by aggregating paragraph-level interactions of case pairs. SAILER(Li et al., 2023a) aims to maximize the utilization of information in annotated data. Specifically, it utilizes an encoder-decoder architecture for pre-training, which encodes the facts into vectors and uses a decoder to re-construct the vectors to the original decisions and reasoning. In addition, Wei et al. (2022) proposes IOT-Match, which learns to find rationales from paired legal cases based on semantics and legal elements of sentences, with the assistance of numerous expert-annotated interpretable relevance labels. Despite the promising results of these models on specific test sets, they still rely heavily on high-quality annotated data. Owing to the scarcity of annotated data, a notable concern arises regarding the limited generalizability of these models across diverse data.
Figure 1. An example of a challenge of legal relevance judgments.
@@
These concerns highlight the role of a more reliable and valid method to collect interpretable legal relevance judgments, which is vital not only for the evaluation of models but also for the training of more sophisticated models. However, to judge relevance, one necessitates substantial domain expertise and a significant effort to comprehend lengthy texts (Shao et al., 2023). This means that manual judgments require much more complex preliminary instructions for annotators, leading to higher learning effort, greater demands on their abilities, and difficulties in expanding the scale of annotated data. In particular, the major challenges in legal cases relevance judgments are threefold: 1. Expertise-intensive: Accurate judgments in the legal domain demands high-level domain knowledge. This includes understanding the intrinsic connection between factual scenarios and their corresponding judicial interpretations. It also requires recognizing how objective conduct and subjective attitudes interact to shape the final legal conclusion. 2. Lengthy-text: The query and candidates often contain thousands of words. 3. Nuance-sensitive: As shown in Figure 1, besides a comprehensive understanding of the overall context, sensitivity to nuances in context and terms is equally important. Because such minor distinctions can significantly impact judicial interpretations and decisions, e.g., ”homicide” vs ”manslaughter”, ”minor injury” vs ”minor assault”, and ”robbery” vs ”Theft”. (Please refer to Section  3.1 Relevance part for further details of legal relevance).
@@
Existing work(Thomas et al., 2023; Faggioli et al., 2023) has shown that the judgment capabilities of large language models (LLMs) in general NLP tasks can rival those of crowd workers. Most efforts often involved general relevance or direct relevance signals, like question-answering, and text classification(Thalken et al., 2023; Gilardi et al., 2023). Since relevance labels in legal cases are far more complicated as mentioned earlier, our major research question is whether LLM can perform relevance judgments in the legal field. Although researchers have developed legal domain-specific large language models (LLMs) such as ChatLaw(Cui et al., 2023) and LawGPT(Hongcheng et al., 2023), these models are primarily fine-tuned for general legal question-answering tasks. Due to limitations in model parameters and the volume of training data, their understanding and reasoning capabilities are inferior to those of general-purpose LLMs like ChatGPT, making it challenging to directly employ them for relevance judgments in long-form legal cases. Furthermore, ChatLaw is trained mainly on English legal data due to the scarcity of high-quality annotated Chinese legal data(Cui et al., 2023). This underscores the critical need to develop and curate more high-quality Chinese legal datasets to enhance the performance and applicability of legal models in Chinese legal contexts.
@@
In this work, we will focus on developing an automatic relevance judgment method for legal case retrieval task. Instead of utilizing or building specialized LLMs, which carry many serious constraints in cost and capabilities (Please refer to the next section for details), we employ one of the most popular general LLMs, ChatGPT, for domain-specific judgments. To take advantage of the understanding and reasoning capabilities of the general LLM, we instruct it step-by-step(Wei et al., 2022) with minimal expert guidance. As a result, our approach can enable a general LLM to make judgments on complex legal case relevance in an interpretable manner, closely aligned with those of human experts. Specifically, we address the challenge of Lengthy-text by decomposing the judgment into fine-grained sub-processes, including distinct views in fact extractions and relevance judgments. In each sub-process, we elaborate expert’s reasoning process as instructions to ease the Expertise-Intensive challenge and mitigate the Nuance-Sensitivity issue. In addition, we employ adaptive retrieval of effective demonstrations to further tackle the challenge of Expertise-Intensive. Given a group of unlabeled legal cases, we also designed a strategy for efficiently collecting possible positive case pairs.
Furthermore, we leveraged label-only synthetic data to fine-tune downstream retrieval models, significantly enhancing the performance of legal case retrieval. Notably, by using synthetic data with both labels and interpretable reasoning generated during the judgment process, we trained two 7B-scale LLMs that match or even outperform GPT-3.5 in judging legal case relevance. These two applications yielded substantial performance gains, providing strong empirical support for the effectiveness of our approach.
@@
  • We design an automatic workflow to leverage a general LLM to make relevance judgments for legal cases. To the best of our knowledge, this is the first attempt at automated legal case relevance judgments.
  • We evaluate and analyze the quality of automatic relevance judgments by comparing them with human labels. Empirical experiments demonstrate that our approach can achieve high consistency with human experts.
  • We use the proposed method to generate a synthetic dataset and show that with minimal expert supervision, our approach enables a large language model to acquire case analysis expertise, and subsequently transfer this ability to a smaller model via annotation-based knowledge distillation.
@@ 2.1. Pre-trained Language Models (PLMs) in Legal case retrieval
Pre-trained Language Models (PLMs) like BERT(Devlin et al., 2018), GPT-2(Lagler et al., 2013), and RoBERTa(Liu et al., 2019) have set new standards in a variety of NLP tasks. The strength of PLMs lies in their ability to understand context, making them effective for tasks ranging from sentiment analysis to machine translation. However, in the legal domain such as legal case retrieval, general PLMs still struggle with the complex nuances and intensive background domain knowledge of legal jargon and may produce biased or asymmetric predictions. Additionally, legal documents typically consist of thousands of tokens, far exceeding the length that mainstream PLMs can handle(Xiao et al., 2021). Researchers are dedicated to designing specialized models(Shao et al., 2020; Li et al., 2023a; Chalkidis et al., 2020; Xiao et al., 2021) tailored to legal case matching, underscoring the limitations of generic PLMs in accurately interpreting and classifying extensive legal content.
@@
Large-scale Language Models (LLMs) are an extension of the PLM concept but on a much grander scale, such as GPT4(OpenAI, 2023) and LLaMa(Touvron et al., 2023a). With billions or even trillions of parameters, LLMs are designed to understand and generate human-like text with little to no fine-tuning required for specific tasks, showing significant potential in various fields, including law.
@@
Although researchers have attempted to train domain-specific LLMs for the legal field(Cui et al., 2023; Hongcheng et al., 2023), there are still several drawbacks in directly training a dedicated large language model for this domain: 1. The process requires a substantial amount of high-quality data and expensive computational resources. Legal texts can be vast and diverse, necessitating extensive training data to ensure the model’s proficiency in understanding the nuances of legal language and reasoning. 2. The legal domain encompasses numerous downstream tasks, each with its own specific requirements and nuances. As a result, domain-specific large-scale language models lack generalizability and reliability across various downstream tasks. Adapting these models to each specific task requires fine-tuning with task-specific data, leading to considerable efforts for collecting specialized training data for each task. 3. Compared to commercial general large models, these domain-specific large models suffer from insufficient training scale and data volume, leading to weaker model memorization, comprehension, and reasoning capabilities. Additionally, most of these domain-specific models have mainly undergone fine-tuning for legal question-answering tasks.
@@
In this article, we present a lightweight approach that enables general LLMs for automated legal relevance judgments.
2.2. LLM Agents for Data Augmentation
@@
Using LLM agents for data augmentation is a promising avenue, especially for tasks where data collection is challenging. (Wang et al., 2022) proposed a method that combines semantic data augmentation with distance metric learning for domain generalization. (Meyer et al., 2022) explored the feasibility of GPT-3 generated synthetic data for training conversational AI classifiers, revealing that while synthetic data offers advantages, it doesn’t surpass real user data in performance. In addition, (Saad-Falcon et al., 2023) develop a method for using large language models (LLMs) to generate large numbers of synthetic queries cheaply, which generates a small number of synthetic queries using an expensive LLM and then uses a much less expensive one to create large numbers of synthetic queries that are used to fine-tune re-rankers. These studies underscore the growing importance and potential of utilizing LLM to generate synthetic data, especially in scenarios with limited data resources. While it offers a novel way to leverage the capabilities of large-scale language models, care must be taken to ensure that the generated data aligns well with real-world scenarios and doesn’t introduce biases or errors. As with any data augmentation technique, it’s crucial to validate its effectiveness empirically on the target task. Therefore, in this work, we will evaluate our automatic judgment approach from two levels: the consistency with human labeling and the empirical effectiveness of data-augmented models on the LCR task.
@@ 3. Methodology
Figure 2. The proposed workflow of data judgments for legal case retrieval task. To be clear, please note that Adaptive Demo-Matching (ADM) is applied twice before Fact Extraction (FE) and Fact Annotation (FA).
@@ 3.1. Preliminary
Relevance: According to LeCaRD(Ma et al., 2021), the Chinese legal case retrieval dataset that we use, of which the data format is shown in Figure 3, the determination of case relevance primarily depends on the key facts of the case, which are composed of ”Material Facts (MF)” and ”Legal Facts (LF)”. Legal Facts are the legal evaluations of Material Facts. For example, in a hypothetical case:
Facts: A had persistent conflicts with B. Consequently, A found a pretext to provoke B and then assaulted him. B’s injuries were classified as ’Level 2 minor injury.
@@
To determine whether different cases are related, one should deep into the Material Facts and the Legal Facts. The Material Facts of a case include the identity of multiple factors such as the defendant and the victim, the cause, process, outcome, time, and location of the crime committed by the defendant. Judging the relevance of MF involves a ”factual level” assessment of the relatedness of different cases, focusing on whether there is ”identity” between multiple factors. Judging the relevance of LF includes the subjective motive of the defendant, whether the victim is at fault, and whether there is a causal relationship between the criminal act and the damage caused. This involves an adaptive ”legal level” assessment in different legal aspects, focusing on whether there is ”homogeneity” between two cases. In practice, the judgment should follow the order of ”Material Facts first, Legal Facts second.” Both are necessary and sufficient conditions for the assessment of case-relevance.
@@
3.2. Automated Relevance Judgments for Legal Cases
@@
We have meticulously designed a prompt workflow for legal case retrieval, which can be segmented into four steps: 1) Preliminary factual analysis by legal professionals, 2) Adaptive Demo-Matching (ADM), 3) Fact Extraction (FE), and 4) Fact Annotation (FA). Figure 2 shows the overall judgment workflow. In the subsequent sections, we will delve into each of these components in detail.
@@
Preliminary factual analysis by legal professionals: While the relevance judgments of the original data were made based on the criteria mentioned in Section 3.1, the annotated data only contain relevance scores without explicit matching labels, preventing the understanding of the exact facts that determine the similarity between two cases. Accurate and direct indications of what the ’Material Fact’ in the current case is, its connection to the details in the text, followed by additional legal knowledge and reasoning logic used for identifying such connections, can motivate LLMs to utilize their inherent knowledge and comprehension abilities to mimic human experts’ cognitive processes. Such an approach is deemed necessary because we observed that, without detailed relevance indications, the model’s judgments tend to deviate and exhibit less explanation consistency. For instance, if case 1 involves person A stealing a car and case 2 involves person B stealing a wallet, the LLM sometimes perceives a car and a wallet as dissimilar items, focusing on trivial facts, thereby simply judging the two cases as irrelevant. Thus, we consider such legal reasoning order as a logical chain, instructing the LLM to make judgments consistent with those of the professionals. Figure 4 and Figure 5 show examples of experts’ demonstrations in FE for MF and LF.
Figure 4. An example of demonstration used in Fact Extraction (FE) for extracting Material Facts (MF). (Translation of Chinese texts.)
Figure 5. An example of demonstration used in Fact Extraction (FE) for extracting Legal Facts (LF). (Translation of Chinese texts.)
Figure 6. Two examples of demonstration used in Fact Annotation (FA) for judging Material Facts (MF) and Legal Facts (LF) respectively. (Translation of Chinese texts.)
@@
As noted in Section 3.1, legal judgments should follow a sequential structure: Material Facts are established first, followed by Legal Facts. Consequently, the Fact Extraction of LF involves a multi-dimensional analysis based on the Fact Extraction of MF, including considerations such as the defendant’s subjective intent, possible contributory fault of the victim, and the causal relationship between the criminal act and the resulting harm. Then, during the Fact Annotation phase, it is crucial to perform adaptive relevance assessments tailored to different case pairs, based on the connection between MF and LF. We incorporate carefully designed expert reasoning logic for relevance analysis to guide the LLM in following experts’ judging criteria.
Formally, for a given stage $t\in\{\text{FE},\text{FA}\}$ and fact type $f\in\{\text{LF},\text{MF}\}$, the expert-curated demonstration sets is denoted as $\mathcal{D}_{t}^{f}$ for a given stage $t\in\{\text{FE},\text{FA}\}$ and fact type $f\in\{\text{LF},\text{MF}\}$, where
[…]
Figure 6 shows an example of demonstrations for the Fact Annotation.
Adaptive Demo-Matching (ADM): Adaptive selection of demonstrations for few-shot in-context learning ensures better alignment between the input and expert instruction, thereby enhancing the LLM’s ability to emulate expert legal reasoning.
Therefore, we use a retrieval function $R(\cdot)$ to adaptively retrieve case-specific demonstrations, which will be placed into the input prompt template. Specifically, given an input case $x$, we retrieve the top-$k$ most relevant examples from the expert-curated demonstration sets. For a given stage $t\in\{\text{FE},\text{FA}\}$ and fact type $f\in\{\text{LF},\text{MF}\}$, we select demonstrations as follows:
$\text{ADM}(x;t,f)=R\left(x;\mathcal{D}_{t}^{f}\right)$
where $R(x;\mathcal{D})$ retrieves the top-$k$ most similar examples to $x$ from $\mathcal{D}$, using BM25 in our setting.
@@
Fact Annotation (FA): As illustrated in Figure 8, we present an framework of FA prompting. Similarly, we separately annotate MF and LF sequentially. Given as input the facts of a pair of cases to be labeled, the prompt contains relevant as well as irrelevant demonstrations. Ultimately, the LLMs are instructed to assess the relevance of the current input pair, drawing upon the expert’s rationale. The calculation of the final relevance score follows the method of LeCaRD: score 1 for MF relevance, and score 2 for LF relevance, with the total relevance score calculated as the sum of MF and LF relevance scores.
Figure 8. The example of Fact Annotation (FA). (Translation of Chinese texts.)